OpenStudy (anonymous):

I need help integrating (1-4x^2)^(1/2)

OpenStudy (amistre64):

it looks almost set up for a sin substitution

OpenStudy (amistre64):

\[\int \sqrt{1-4x^2}dx\] \[\int \sqrt{4(\frac{1}{4}-x^2)}dx\] \[2\int \sqrt{(\frac{1}{4}-x^2)}dx\] \[2\int \sqrt{(\frac{1}{4}(1-sin^2(t))}\ \frac{1}{2}(cos(t))dt\] \[2\int \frac{1}{2}cos(t)\frac{1}{2}cos(t)dt\] \[\frac{1}{2}\int cos^2(t)dt\]

OpenStudy (amistre64):

\[\frac{1}{2}\int cos^2(t)dt\] \[\frac{1}{2}(\frac{cos(t)sin(t)}{2}+\frac{t}{2})\] cos(t) = \(\sqrt{1-4x^2}\); sin(t) = 2x ; t=\(sin^{-1}(2x)\) \[\frac{1}{2}(\frac{(\sqrt{1-4x^2})(2x)}{2}+\frac{sin^{-1}(2x)}{2})\] \[\frac{1}{2}(\frac{(\sqrt{1-4x^2})(x)}{1}+\frac{sin^{-1}(2x)}{2})\] \[\frac{x\sqrt{1-4x^2}}{2}+\frac{sin^{-1}(2x)}{4}+C\]